OpenStudy (ttop0816):

Simplify 1 · x - x/1. 0 1 x

OpenStudy (anonymous):

Is the expression: \[1\times\frac{x-x}{1}\] or \[1\times x-\frac{x}{1}\]

OpenStudy (ttop0816):

so would it be 1?

OpenStudy (anonymous):

Well is it the top expression or the bottom expression. Im kinda confused as to the question...

OpenStudy (anonymous):

Although I suppose either way...it would be zero...

OpenStudy (ttop0816):

hmmm.... okay! thanks for your help though! (: